2.14x - 12.18 = -5.76
Answer by checkley77(12844)

2.14x - 12.18 = -5.76
2.14X=-5.76+12.18
2.14X=6.42
X=6.42/2.14
X=3 ANS.
PROOF:
2.14*3-12.18=-5.76
6.42-12.18=-5.76
-5.76=-5.76